### v2

# Time arrow in open-boundary one-dimensional stochastic dynamics

Abstract
We consider the finite-timestep Brownian dynamics of a single particle confined in one dimension, with a nonuniform temperature profile. In such an open-boundary scenario, one cannot observe any net probability current in the nonequilibrium steady state (NESS). On the other hand, the nonequilibrium nature of this system is exhibited through the asymmetry in forward and backward transition probabilities, as is reported in this work through the stochastic simulation analysis and theoretical arguments. The irreversibility becomes prominent nearby the temperature interface. We propose that the observed irreversibility can be accounted for via a virtual-gyration scenario, while the collapse of virtual gyrations upon the one-dimensional coordinate leads to the absence of probability current.
I Introduction
In the discussions of stochastic dynamics, the systems of autonomous gyrators[1, 2, 3, 4, 5, 6, 7] have recently drawn great interest. For these systems, the most renowned feature is the emergence of a perpetual, average circulating motion, despite the fact that the dynamics is highly random within short-time intervals. In two-dimensional stochastic systems, such gyrating behavior can arise due to the asymmetry in thermal fluctuations along the two coordinates[1, 5, 6], or due to the temperature inhomogeneity in space[9, 8]. The understanding of such autonomous gyrating systems is crucial for the development of microscopic heat engines.
In one-dimensional systems, persistent average gyration has been studied since the famous Büttiker-Landauer discussions in the 1980s[10, 11]. Through the closed-loop topology, a nonvanishing probability current can be achieved through a proper setting of temperature and potential profiles. The directed average dynamics, as an indicator of time irreversibility, is an important characteristic of the second law in thermodynamics, or its recent variants in stochastic dynamics, the fluctuation theorems[12, 13, 14, 15, 16]. In the Büttiker-Landauer engine, as well as in two-dimensional Brownian gyrators, it can be shown that the average net heat dissipation along the gyrating loop is positive. This is in agreement with the second law in the sense that the overall change in total entropy along a closed autonomous cycle is, on average, positive. Through the successful development of fluctuation theorems, scientists proceed to rephrase the second law of thermodynamics with more probability usage.
If one considers the one-dimensional stochastic dynamics, just like the system of a Büttiker-Landauer engine, but with an open-boundary condition instead, the steady-state dynamics appears trivial with a zero average probability current over all positions. On the other hand, regarding thermodynamics, the system must stay in nonequilibrium under a nonuniform temperature profile. Without a persistent gyrating current, one may be tempted to search about the signature of time irreversibility for such a system. In this current work, we study the one-dimensional open-boundary system with discrete transition timesteps. We report that time irreversibility can indeed be observed, and virtual “gyrations” can occur on the local scale, for transitions near the temperature discontinuity.
The nonequilibrium nature in similar temperature-gradient systems has been discussed in Refs. [8, 17, 18, 19]. These studies revealed the discrepancy in overdamped stochastic dynamics, as well as its shortage of contribution in total entropy dissipation of temperature-gradient systems. For example, in the work of Celani et al.[18], they obtained in underdamped stochastic dynamics an additional source of time irreversibility, or “entropy anomaly” in overall dissipation. This irreversible behavior is attributed to the circulating steady-state behavior in the phase space, and such an entropy anomaly persists even if one proceeds towards the overdamped limit. In our current work, we consider the overdamped stochastic dynamics only, and we find yet another source of time irreversibility, which can arise due to the finiteness of time intervals between successive thermal kicks. And in contrast to the work in [18], which assumes a smooth temperature variation over space, we choose to study the other extreme, in which temperature changes abruptly.
Our work is structured as follows: in Sec. II, we introduce our system of study, its dynamic stochastic equation, as well as the corresponding Fokker-Planck equation analysis that works for infinitesimal timesteps. In Sec. III, we compute the finite-timestep transition probabilities and demonstrate through a crude approximation that time reversibility, or its synonym, detailed balance, may be violated. We go on to examine the violation of detailed balance through theoretical arguments. In Sec. IV we present our simulation result that confirms the existence of time irreversibility, and meanwhile, the steady-state probability distribution obtained from simulations exhibits deviation from the Fokker-Planck equation analysis. In Sec. V, we seek to fit the result of the probability distribution to that of a corresponding continuous-timestep model, through the introduction of an effective temperature profile. We remark that the smoothed-out characteristic of effective temperature can be interpreted by an averaging effect of the temperature profile over a single diffusion timestep.
II System
We consider a Brownian particle moving in one dimension, its movement being dictated by the Langevin equation
$\gamma\frac{dx}{dt}=-\frac{dU}{dx}+\sqrt{2\gamma k_{\rm B}T(x)}\,\xi\,.$ (1)
Thus, a Brownian particle of position $x$ is subject to a smooth potential $U(x)$, while the last term in the equation above gives rise to the random force. This fluctuating force is Gaussian white and uncorrelated, namely, $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$. For simplicity, all the physical quantities we use in this work are set to be dimensionless. Under this representation, the Boltzmann constant and $\gamma$ are both assigned to be unity. The latter can be achieved through a rescaling of time by the factor $1/\gamma$. The temperature is profiled by $T(x)=T_{H}$ for $x>x_{0}$ and $T(x)=T_{C}$ for $x\leq x_{0}$. Note that the product $\sqrt{T(x)}\,\xi$ in Eq. (1) adopts Itô’s scheme.
The stochastic dynamics of Eq. (1), under the dimensionless representation, can be equivalently described via the Fokker-Planck equation (FPE):
$\frac{\partial P}{\partial t}=\frac{\partial}{\partial x}\left[\frac{\partial(TP)}{\partial x}+P\frac{\partial U}{\partial x}\right]\,,$ (2)
where $P$ is the probability density of the particle. Note that the equivalency between the descriptions in Eqs. (1) and (2) are based on the scenario when random kicks are continuously dispersed over time. In the nonequilibrium steady-state (NESS) regime, one then has $\frac{d\,j}{dx}=0$, while
$j=-\left[\frac{\partial(TP)}{\partial x}+P\frac{\partial U}{\partial x}\right]$ (3)
represents the probability current. Upon reaching the steady state, $j=\rm{constant}=0$ everywhere in our one-dimensional open-boundary system. Therefore, within each constant-temperature domain, the probability density is proportional to the Boltzmann distribution. To determine the coefficients for the probability densities of these two regions, we consider an infinitesimal transition across the temperature discontinuity $x=x_{0}$. Since Eq. (3) implies $d(TP)+PdU=0$ in the NESS, if $PdU$ is negligible for infinitesimal transition step, so is $d(TP)=0$. Therefore, one has
$(TP)|_{x=x_{0}^{-}}=(TP)|_{x=x_{0}^{+}}\,.$ (4)
Note that the discontinuous temperature profile leads to a probability density discontinuity. Thus, according to the Fokker-Planck equation, one has
$P={\rm{constant}}\cdot\frac{e^{-\beta(U-U_{0})}}{T}\,,$ (5)
where $\beta\equiv 1/T$, and $U_{0}$ is the potential energy right at the temperature discontinuity.
Note that the FPE represents a special case of the master equation, following the truncation of the Kramers-Moyal expansion[20, 21] up to the second order. The legitimacy of truncation is based on the underlying assumption that transitions are local. In particular, for the one-dimensional case, the FPE can faithfully represent the statistics of random walks over nodes with a chain-like connecting structure as illustrated in Fig. 1(a). Regarding this open-chain structure, its steady-state distribution meets the detailed-balance criterion if the probability vanishes at far distances. However, in the following sections, we demonstrate that the detailed-balance condition, and hence time reversibility, fail in this model if random kicks occur over discrete timesteps and the temperature profile exhibits discontinuity. In such a case, the open-chain structure cannot represent the system faithfully. Meanwhile, this failure also leads to a notable deviation from the FPE analysis.
Figure 1: Illustrations of two scenarios. Arrows represent probability currents. Points $C$, $B$, $A$ represent successive discrete positions from left to right, while point $A$ is at a higher temperature. (a) The chain-connecting scenario. Note that the net steady state current between nodes has to be zero everywhere (detailed balance). (b) The loop scenario. In finite-timestep stochastic dynamics, point $A$ can reach $C$ directly over one single timestep $\Delta t$, thanks to the large diffusion parameter in $A$. This feature gives rise to the loop-like scenario and can contribute to a virtually gyrating current at the steady state.
III Transition Probability; Violation of Detailed Balance
Now, let us study the Brownian motion again. This time, the random kicks emerge over discrete timesteps. Consider a transition from position $x$ to $x^{\prime}$ over timestep $\Delta t$. The Langevin equation then reads
$x^{\prime}-x=F(x)\Delta t+\sqrt{2T(x)}\,\xi\,\Delta t\,,$ (6)
where $F=-\frac{dU}{dx}$ represents the conservative force, and $\xi\,\Delta t$ corresponds to the noise of a typical Wiener process. From Eq. (6), one derives the conditional transition probability
$P(x^{\prime}|x)=\frac{1}{\sqrt{2\pi}}\cdot\frac{1}{\sqrt{2T(x)\,\Delta t}}\exp\left[-\frac{\Delta t}{4T(x)}(\dot{x}-F(x))^{2}\right],$ (7)
and that for the backward transition (i.e., $x^{\prime}\to x$):
$P(x|x^{\prime})=\frac{1}{\sqrt{2\pi}}\cdot\frac{1}{\sqrt{2T(x^{\prime})\,\Delta t}}\exp\left[-\frac{\Delta t}{4T(x^{\prime})}(-\dot{x}+F(x^{\prime}))^{2}\right].$ (8)
Let us use the abbreviation that $T^{\prime}\equiv T(x^{\prime})$ and $F^{\prime}\equiv F(x^{\prime})$, while the unprimed letters $T$ and $F$ are reserved for the quantities at position $x$. Comparison of the forward and backward conditional transition probabilities then gives
$\ln\frac{P(x^{\prime}|x)}{P(x|x^{\prime})}$ $=\frac{1}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left[-\frac{(\dot{x}-F)^{2}}{T}+\frac{(\dot{x}-F^{\prime})^{2}}{T^{\prime}}\right]$
$\approx\frac{1}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left[\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(\dot{x}^{2}+F^{2})+2\dot{x}F\left(\frac{1}{T}+\frac{1}{T^{\prime}}\right)\right]\,.$ (9)
Note that the approximation of $F^{\prime}$ by $F$ results in errors of magnitude $O(\Delta x^{2})$ and $O(\Delta t\cdot\Delta x)$ in the last two terms in Eq. (9), respectively ($\Delta x\equiv x^{\prime}-x$). For a constant-temperature transition $T^{\prime}=T$, Eq. (9) reduces to the heat entropy dissipation $(F\cdot\Delta x)/T$[16]. Meanwhile, regarding a transition crossing over different temperature regions, Eq. 9 reveals additional sources of entropy dissipation due to the temperature change.
To get a quick look at time irreversibility, let us assume for the time being that the solution of FPE is valid. With the use of Eq. (5) and (9), one then has
$\ln\frac{P(x^{\prime},x)}{P(x,x^{\prime})}$ $=\ln\left[\frac{P(x)}{P(x^{\prime})}\cdot\frac{P(x^{\prime}|x)}{P(x|x^{\prime})}\right]$
$\approx\frac{3}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(\dot{x}^{2}+F^{2})-\frac{\Delta U}{2}\left(\frac{1}{T}+\frac{1}{T^{\prime}}\right)-\frac{1}{T}(U-U_{0})+\frac{1}{T^{\prime}}(U^{\prime}-U_{0})$
$=\frac{3}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(\dot{x}^{2}+F^{2})+\frac{1}{2}\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(U+U^{\prime}-2U_{0})\,,$ (10)
where $P(x^{\prime},x)$ represents the joint probability that the random walker is at site $x$ at time $t$ and $x^{\prime}$ at time $t+\Delta t$. Therefore, the detailed-balance condition holds for a uniform-temperature profile, at least to the order of magnitude considered here. On the other hand, Eq. (10) predicts time irreversibility for transitions across a temperature discontinuity.
Next, we proceed to show again the existence of time-irreversibility, or equivalently, detailed-balance violation, this time without assuming the validity of Eq. (5). Let us consider three neighboring sites $A,B$ and $C$, while the temperature at site $A$ is $T_{H}$ and the other two sites at a lower temperature $T_{C}$. The distances between $AB$ and $BC$ are identically equal, and we denote this distance to be $\Delta x$. Let us first consider the identity
$\frac{P(A,B)P(B,C)P(C,A)}{P(B,A)P(C,B)P(A,C)}=\frac{P(A|B)P(B|C)P(C|A)}{P(B|A)P(C|B)P(A|C)}\,.$ (11)
With Eq. (9), it leads to
$\ln\frac{P(A,B)P(B,C)P(C,A)}{P(B,A)P(C,B)P(A,C)}$ $\approx\frac{(\Delta x)^{2}}{4\,\Delta t}\left(\frac{1}{T_{H}}-\frac{1}{T_{C}}\right)+0+\frac{(2\Delta x)^{2}}{4\,\Delta t}\left(\frac{1}{T_{C}}-\frac{1}{T_{H}}\right)$
$\ \ +\frac{\Delta x}{2}\left(\frac{F}{T_{H}}+\frac{F}{T_{C}}\right)+\frac{\Delta x}{2}\left(\frac{F}{T_{C}}+\frac{F}{T_{C}}\right)-\frac{2\Delta x}{2}\left(\frac{F}{T_{H}}+\frac{F}{T_{C}}\right)$
$=\frac{3}{4}\frac{(\Delta x)^{2}}{\Delta t}\left(\frac{1}{T_{C}}-\frac{1}{T_{H}}\right)+\frac{F\,\Delta x}{2}\left(\frac{1}{T_{C}}-\frac{1}{T_{H}}\right)\,.$ (12)
In our work we choose $(\Delta x)^{2}=O(\Delta t)$. Therefore, the magnitude of $\Delta x$ is equal to the diffusive scale over time $\Delta t$ if $O(T_{H})=O(T_{C})=1$. In this case, the first term in the final line of Eq. (12) outweighs the second term, and one observes significant violation in the detailed-balance condition: at least one forward-to-backward ratio of the $(A,B)$, $(B,C)$, $(C,A)$ pairs is not equal to 1.
Finally, we argue that the detailed-balance violation can occur for transitions over constant-temperature sites. Assuming that for our system of consideration, the detailed-balance condition still holds for constant-temperature transitions, then there would be no probability current between constant-temperature nodes. Let us expand our nodes of consideration such that the new node $A^{\prime}$ consists of all high-temperature sites, $B^{\prime}$ represents site $B$ only, while $C^{\prime}$ consists of all low-temperature sites other than site $B$. Thus, the assumption of detailed balance between $B^{\prime}$ and $C^{\prime}$ leads to the conclusion of no probability current between them. Yet a 3-node steady state cannot be achieved in this manner, if there exist nonzero currents between $(A^{\prime},B^{\prime})$ and/or $(A^{\prime},C^{\prime})$ (see Fig. 1(b)). Therefore,, we come to the conclusion that the detailed balance can be violated even between constant-temperature nodes. And in turn, it implies the failure of Eqs. (10) for constant-temperature transitions. This failure originates from the deviation in steady-state probability distribution from Eq. (5).
IV Simulation results
In this work, we consider the harmonic potential $U=x^{2}/4$. The temperature discontinuity is set to the position $x_{0}=-2$, while we employ $T_{H}=4$ and $T_{C}=2$. Our numerical result was obtained via Eq. (1) using the discrete timestep $\Delta t=10^{-2}$. The simulation was performed over a total of $10^{8}$ timesteps to achieve sufficient sampling of the NESS behavior. The recorded time series in position is classified into bins of width $2\delta$, so that statistical distributions can be obtained.
IV.1 Joint Transition Probability
To quantitatively assess the violation of detailed balance predicted in section III, we compare the asymmetry in forward and backward transition probabilities across multiple spatial points. Specifically, we look at the joint probability $P(x^{\prime},x)$ of finding the particle at position $x$ at time $t$ and position $x^{\prime}$ at a later time $t+\Delta t$, and compare it with the occurrence of the backward process $P(x,x^{\prime})$. The joint probabilities are obtained via trajectory analysis. Note that due to the coarse-graining procedure, the notions $x$ and $x^{\prime}$ in $P(x^{\prime},x)$ in our data analysis actually mean all possible points residing in the bins centered at $x$ and $x^{\prime}$, respectively.
We first divide the spatial domain into small intervals of width $2\delta$, and compute the ratio of forward to backward transitions between adjacent intervals:
$R(x)\equiv\frac{P(x+\delta,x-\delta)}{P(x-\delta,x+\delta)}\,.$ (13)
In this work, we choose $\delta=0.1$, which is similar to the diffusing distance over a single timestep $\Delta t$. The result, shown in Fig. 2, indicates significant deviation of $R(x)$ from unity near the temperature discontinuity, signifying the violation of detailed balance. Note that for single-step transitions over identical-temperature regions, Eq. (9) reduces to $-\Delta U/T$. According to this prediction, the ratio of joint probabilities $R(x)$ can be represented as
$R(x)=\frac{P(x-\delta)}{P(x+\delta)}\cdot\exp{(-\Delta U/T)}\,.$ (14)
or equivalently, the ratio of $P(x)\cdot\exp{[U(x)/T]}$ over successive intervals. With the use of NESS probability distribution extracted from the simulation result, the prediction by Eq. (14) matches well with the joint-probability ratio Eq. (13). The only exception occurs at the point $x=x_{0}$, where the forward and backward transitions cross over the temperature discontinuity. In this case, Eq. (14) cannot be applied.
Figure 2: Ratio of joint probabilities between forward and backward transitions. The counting region width $\delta$ is determined by the global root-mean-square (RMS) displacement of the Brownian particle. Blue dots: result where joint probabilities were extracted directly from simulation data; green dots: result evaluated from Eq. (14), via the NESS probability distribution $P(x)$ from simulation.
IV.2 NESS probability distribution
Note that the FPE analysis Eq. (5) results in the Boltzmann-like behavior in NESS probability distribution over constant-temperature intervals. Along with the boundary condition Eq. (4), the FPE analysis gives
[…]
where Z is the normalization constant. If the FPE analysis were legitimate, according to Eq. (14), detailed-balance violation could not be observed in transitions over constant-temperature intervals, as the violation is demonstrated in Fig. 2. The direct comparison between the FPE prediction of the NESS probability distribution and the simulation results, as shown in Fig. 3, also reveals some discrepancy. This deviation is particularly pronounced near the temperature discontinuity at $x=x_{0}$. The discrepancy suggests about the inaccuracy of the FPE prediction in the case of discrete-time transitions.
Figure 3: Comparison of NESS probability distributions $P(x)$ for a Brownian particle in a harmonic potential with inhomogeneous temperature profile. Blue dots represent the simulation results; the orange solid line shows the theoretical prediction from Eq. (15) using the step-function temperature profile; the green dashed line represents the modified theoretical prediction using the effective temperature profile from Eq. (17) (see inset).
IV.3 Study in NESS probability current
In a one-dimensional open system, the NESS probability current $j(x)$ must be zero everywhere. This can be understood via the following argument: if one sets up a monitoring station at any arbitrary position, and count the number of events for a random walker to go through this site (the count number gains “+1” if the particle is passing from left and “-1” vice versa). And over the long course of a simulation, the total count number can have 3 possible values only: +1, 0, and -1, because the random walker cannot pass the station twice from the same side without returning back in between. In our simulation work, we try to numerically examine the NESS probability current via the formula $j(x)=P(x)v(x)$, where $v(x)$ represents the corresponding average velocity at site $x$. The latter is obtained by the finite-difference method
$v(x)\approx\frac{1}{2\Delta t}\langle[(x_{t+\Delta t}-x_{t})+(x_{t}-x_{t-\Delta t})]|x_{t}=x\rangle\,.$ (16)
The result of the computed $j(x)$, as presented in Fig. 4, exhibits noticeable deviation nearby the temperature discontinuity. The deficit is attributed to the fact that the computation scheme fails to count certain events. To be more specific, in the current model of description, the events that are generated from diffusions in the high-temperature region possess higher mobility. And these “fast” events, of which the random walker “bypasses” the monitoring bin over timestep $\Delta t$, often fail to be counted in the calculation of average velocity at the midpoints of their paths using Eq. (16) (see Fig. 1(b)). The deficit in the calculated probability current is thus more prominent nearby the temperature interface, where the calculated average velocity becomes positive.
Figure 4: Probability current $j(x)$, evaluated via $j=P(x)v(x)$, where $v(x)$ is obtained through finite difference approximation of particle velocities. Note that for one-dimensional open-boundary systems, the ideal probability current has to be zero at all positions.
V Modified Temperature Profile and Probability Distribution
The comparison of the simulation result and the FPE prediction in the probability distribution shown in Fig. 3 shows that the simulation result is less discontinuous in slope across the temperature discontinuity. Meanwhile, the discrepancy triggers the following physical consideration: in the discrete-time simulation, what is the representative temperature of a Brownian particle crossing a sharp discontinuity at $x_{0}$. Nearby the temperature discontinuity, the particle that arrives at a position of consideration at time $t+\Delta t$ may come from positions of either $T_{H}$ or $T_{C}$ at time $t$. The above consideration prompts us to adopt a naïve picture with the use of an “effective” temperature, which differs from the actual step-function profile. In fact, any arbitrary NESS probability distribution must satisfy, in the mathematical sense, some one-dimensional Fokker-Planck-type equation with the introduction of an effective temperature profile. By the choice of a proper effective temperature profile, the probability distribution of our discrete-time stochastic dynamics can still be represented by a Fokker-Planck equation. To fit our simulation result, we propose the following effective temperature profile:
$T_{\text{eff}}(x)=T_{C}+(T_{H}-T_{C})\cdot\frac{1}{2}\left[1+\text{erf}\left(\frac{x-x_{0}}{w}\right)\right]\,,$ (17)
with the characteristic width $w$ given by
$w=\sqrt{2(T_{H}-T_{C})\Delta t}\,.$ (18)
The temperature profile that interpolates between $T_{C}$ and $T_{H}$ utilizing the error function seems to serve as an appropriate candidate, as we expect that the smoother transition in temperature can lead to a less sharp behavior in probability density change near the crossing. The width $w$ is related to the difference in the variances of the diffusion steps between the high- and low-temperature intervals. Thus $w$ depends on the temperature difference $(T_{H}-T_{C})$ and the timestep $\Delta t$. Either a larger temperature jump or a longer timestep leads to a broader spread in the effective temperature transition.
In Fig. 3, we find that, with the use of the effective temperature profile, the resulting NESS probability distribution of FPE analysis fits remarkably well with the one obtained by simulation. We have compared the simulation results with the FPE analysis using the effective temperature profile over various timesteps and temperature settings, while both results stay in pretty good agreement with each other.
The above comparison suggests that one can still derive the distribution of a one-dimensional stochastic simulation over discrete timesteps via an effective FPE analysis. However, this interpretation may appear to be paradoxical, since we noted earlier that any one-dimensional Fokker-Planck-type equation implies an underlying chain-connecting structure. And how could the detailed balance be violated at NESS if the underlying structure is chain-like with open boundaries? We speculate that this paradox may be resolved knowing that time irreversibility is implicitly incorporated through the effective temperature formalism. To be more specific, an effective temperature at a particular position contains the information of its last-step temperature, and this memory effect is in fact irreversible.
VI Summary
Through this study, we demonstrate the existence of time irreversibility in one-dimensional open-boundary stochastic dynamics when random thermal kicks occur over finite time intervals. Similar to the discussions in other areas of nonequilibrium steady-state stochastic dynamics, the irreversibility observed can be represented via some probability current “gyrating” among positions nearby the interface of temperature discontinuity (see Fig. 1(b)). The virtual gyration arises due to the mismatched diffusion distances between the high- and low-temperature adjacent regions. In terms of probability, the joint probabilities of forward and backward transitions exhibit differences nearby the temperature discontinuity, while the steady-state probability distribution in position also deviates from the description of the typical FPE analysis.
The characteristic width $w$, within which the aforementioned nonequilibrium phenomena could be observed, would shrink as the timestep between successive thermal kicks becomes smaller. From Eq. (18), one finds that $w^{2}/\Delta t$ bears the same order of magnitude as the diffusion coefficient. Taking the example of colloidal dispersions in water, this means that the time interval between successive kicks has to be in the order of seconds for $w$ to lie within the observation limit. While this is largely deviated from the natural condition for colloidal suspensions, the nonequilibrium facts discussed here may be observed more likely in systems with artificial noises[22, 23, 24, 25, 26, 27]. The realm of stochastic dynamics with discrete random kicks can be commonly observed in both natural and laboratory systems. Yet more extensive studies need to be conducted in order to explore the potentially distinctive behavior of its continuous-time counterpart. Finally, we remark that while we focus on the open-boundary one-dimensional system, the mechanism responsible for the irreversibility remains valid for the closed-loop Büttiker-Landauer engine, or higher-dimensional stochastic systems with temperature discontinuity.
Acknowledgements.
Acknowledgment: This work has been supported by the National Science and Technology Council of Taiwan under grant no. 113-2112-M008-018-MY2 (PYL).
data availability
The data that support the findings of this article can be found in Ref. [28].